OpenStudy (anonymous):

Solve each equation on the interval 0 less than or equal to theta less than 2pi 2sintheta+1=0

OpenStudy (anonymous):

2sintheta = -1 sintheta= -1/2 as sin is negative , therefore the value of theta lies in third quadrant and fourth quadrant

OpenStudy (anonymous):

sintheta = sin (pi + pi/3) or sintheta = sin(2pi- pi/3) theta = 4pi/3 or theta = 5pi/3